A young man pushes down on a plastic cube with a force of 30.0 N. The 8.5 cm high cube is compressed to 5.5 cm. What is the stress on the cube?
1
Expert's answer
2020-05-04T12:41:45-0400
"\sigma=\frac{F}{A}=\frac{30}{0.085^2}=4150\ Pa"